OpenStudy (anonymous):

which of the following to the equation c+(4-3c)-2=0 a. -1 b. 0 c. 1 i pick this one d.2

OpenStudy (anonymous):

C

OpenStudy (anonymous):

you are correct according to me

OpenStudy (anonymous):

yup its right

OpenStudy (dumbcow):

combine like terms -2c +2 = 0 -2c = -2 c = 1 good job

OpenStudy (anonymous):

C

OpenStudy (anonymous):

I got 1 too